OpenStudy (lifeisadangerousgame):

Can someone show me how to answer all of these questions? Given the inequality below. Will the boundary line be dotted or solid? (5 points) Explain your answer. (5 points) Do the values on the boundary line make the inequality true or false? (5 points) Explain your answer. (5 points) 3x - 5y > 15

OpenStudy (anonymous):

It will be dotted as there is no equal sign. False as they are not smaller or larger than 15 but exactly equal

OpenStudy (lifeisadangerousgame):

My teacher won't accept that though, can you show me how to explain it?

OpenStudy (lifeisadangerousgame):

And my comp didn't copy it correct. \[3x - 5y \ge 15\] that's the correct one

OpenStudy (aroub):

It is solid then

OpenStudy (aroub):

Because there is an equal sign

OpenStudy (lifeisadangerousgame):

Thank you, how do I find if the values on the boundary line make the inequality true or false? (5 points

OpenStudy (aroub):

Do you have to graph it ?

OpenStudy (lifeisadangerousgame):

Yea, but I don't know how

OpenStudy (lifeisadangerousgame):

I know that the shaded area will be above the line

OpenStudy (lifeisadangerousgame):

Should I put (0,0) into the equation to figure tihs out?

OpenStudy (aroub):

Yes

OpenStudy (lifeisadangerousgame):

\[3x - 5y \ge 15\] \[-5y \ge -3x + 15\] \[y \le 3/5x - 3\] and finally \[y \le 3/5x - 3 = 0 \le -3? \]

OpenStudy (lifeisadangerousgame):

Is this right?